(58) If the total bill at a restaurant , including a 15% tip, is $65.32, what was the cost of the meal alone?
(38) A survey of 400 people found that 66 were left-handed. What percent of those surveyed were left-handed?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

X+.15X=65.32
1.15X=65.32
X=65.32/1.15
X=56.80 FOR THE MEAL BEFORE THE TAX IS ADDED
PROOF
56.80*1.15=65.32
65.32=65.32
#######################################################
MERRY CHRISTMAS

Answer by stanbon(48558)

58) If the total bill at a restaurant , including a 15% tip, is $65.32, what was the cost of the meal alone?
Let "x" be the cost of the meal.
EQUATION:
x+0.15x=65.32
1.15x=65.32
x=$56.80 (Cost of the meal before the tip was added)
---------------

(38) A survey of 400 people found that 66 were left-handed. What percent of those surveyed were left-handed?
66/400 = 0.165 = 16.5% were left-handed
------------
Cheers,
Stan H.